OpenStudy (anonymous):

Be sure to distribute and collect like terms before you solve. 2(2x - 4) + 3x = -1 solve for X

OpenStudy (anonymous):

distribute 4x - 8 + 3x = -1 combine like terms 7x - 8 = -1 add the 8 7x = 7 x = 1

OpenStudy (anonymous):

2(2x-4)+3x=-1 4x - 8 + 3x = -1 +8 +8 ___________________ 4x+3x=7 7x=7 x=1